• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Yi Qiang

    Log entry time 22:37:05 on November 04, 2008

    Entry number 246743

    keyword=Target activities during controlled access

    1. After opened the target enclosure, we found out the optics target was severely damaged. More details can be found in Brad's early report: http://www.jlab.org/~adaq/halog/html/0811_archive/081104182808.html

    2. Joe prepared the spare target ladder with 6 carbon foils and we replaced the damaged one with the spare one.

    3. Chiranjib and I swapped the two damaged cable from vertical with the one from the unused longitudinal line. The connections can be found in Jin's report: http://www.jlab.org/~adaq/halog/html/0811_archive/081104214604.html

    4. Re-aligned the transverse line, now both p wave and s wave are hitting the center of cell.

    5. Secured the fiber for transverse line spectrum analyzer, no improvement of the absorption spectrum seen though. I also used RTV to seal the possible leaks around the fiber, however, the major source of the leak is actually coming from the bottom piece which cannot be sealed. So the fiber is still kind of hot, but can be touched this time.

    6. Re-connected the GPIB cables from Fluke45 current meters (for holding field coils) to EPICS IOC

    7. Removed the target wall collimators

    8. EPR RF amplifier seems to be cool, so I decided to leave it in the lead shielding.

    A copy of this log entry has been emailed to: jpchen@jlab.org, brads@jlab.org, chiran@jlab.org,jkatich@jlab.org,jinhuang@jlab.org,zhangyi@jlab.org